2 boys xombined age 66 one is 6 years older, how old are they
Answer by jim_thompson5910(35256)

x+y = 66
y = x+6


x+y = 66

x+x+6 = 66

2x+6 = 66

2x = 66-6

2x = 60

x = 60/2

x = 30

y = x+6

y = 30+6

y = 36


So the two ages are 30 and 36


-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------